Three years ago, Tom was three times as old as Paul will be in seven years time. If Paul is 20 years of age now, how old is Tom?

Answer by lwsshak3(11628)

Three years ago, Tom was three times as old as Paul will be in seven years time. If Paul is 20 years of age now, how old is Tom?
**
let x=Tom's present age
3 yrs ago:
Tom's age=x-3
..
7 yrs hence
Paul's age=20+7=27
..
x-3=3*27
x=81+3=84
Tom's present age=84